OpenStudy (anonymous):

Solve for x by completing the square: -2x^2+2x+10=0

OpenStudy (anonymous):

\[\LARGE x_{1/2}={-b\pm\sqrt{b^2-4ac}\over 2a}\] \[\LARGE ax^2+bx+c=0\] in your case you have: \[\LARGE \underbrace{-2}_{a}\cdot x^2\;\;\;\;\underbrace{+2}_{b}\cdot x\;\;\;\; \underbrace{+10}_{c}=0\]

OpenStudy (anonymous):

now substitute...

OpenStudy (anonymous):

-2x^2+2x+10=0 ---> devide both sides by -2 x^2 - x - 5 = 0 ----> transpose 5 to other side. x^2 -x = 5 ----> divide the only x term by 2 and square it. x^2 - (x/2)^2 = 5 -----> know add the answer you got from the step above to both sides which is 1/4. x^2 + x + 1/4 = 5 + 1/4 ----> the square is competed just simplify it. (x+1/2)^2 = 21/4 ----> simplify again. x + 1/2 = (sqrt(7))/2 x = [(sqrt(7))/2] - 1/2 x = ([sqrt(7)] - 1)/2